OpenStudy (hoopcoach99):

Write an equation in slope-intercept form of the line with the given slope and point: slope = 3 and (1, -2). A. y = 3x - 5 B. y = 3x - 2 C. y = 3x + 1 D. y = 3x + 3

OpenStudy (anonymous):

y = mx + b slope(m) = 3 (1,-2) x = 1 and y = -2 now we sub and solve for b -2 = 3(1) + b -2 = 3 + b -2 - 3 = b -5 = b so your equation is : y = 3x - 5 any questions ?